OpenStudy (anonymous):

Solve for y. 2x + 3y = 3

OpenStudy (anonymous):

subtract 2x from both sides

OpenStudy (anonymous):

then divide everything by 3

OpenStudy (anonymous):

2x + 3y = 3 3y = 3 - 2x y = (3 - 2x)/3